Sadie White receives an hourly rate of $30, with time and a half for all hours worked in excess of 40 during a week. Payroll data for the current week are as follows: hours worked, 48; federal income tax withheld, $300; social security tax rate, 6.0%; and Medicare tax rate, 1.5%. What is the net amount to be paid to White?

$1,143.00

$1,260.00

$1,443.00

$1,000.00
Answer #1

Net Amount to be paid to white

Wages for the current week

For 40hours =40*30                                =1200

For excess of 40(48-40=8)=8*45(30+15)=360

Total                                                         =1,560

Less:Federal Income tax withheld           =300

Less:Social security tax(6%*1560)          =93.60

Less:-Medicare tax rate(1.5%*1,560)      =23.40

Amount to be paid to white                      =$1,143

Answer is A